Enter the correct number to make this a valid probability mass function. 3 10 P(X = x) 10.12 0.1
Answer #1

The probability mars function x 1 3 10 PCX=x) 0.12 0.1 ? EP (G=x;) = P(x=1) + P(2=3) + P(3=10) = 1 0-12 tool + P(x=10) = 1 P(